I am trying to help my son and I am lost. He says his answer must be in a fraction but I can even remember how to work these 2 problems. I think this is how I would write it
#1 (3x+2)/5=7/10

#2 (7-3x)/4=-x/8

Please help

Answer by ikdeep(189)

(3x+2)/5=7/10
multiply both sides by 5 and we get
3x+2 =7/2
subtract 2 from both sides and we get
3x = (7/2)-2
on solving RHS we get ...
3x = 3/2
now divide both sides by 3 and we get
x = 1/2
Now if you want to verify you answer ,,,you can put the values of x in the equation and if you get LHS=RHS,,,this means that you answer is correct..
Try to solve the other problem in the same ways ,,,,
Still if you won't be able to solve further, contact me again.
hope this will help you
Please feel free to revert back for any further queries .